# Pareto-Optimal Estimation and Policy Learning on Short-term and Long-term Treatment Effects

### 05 Mar 2024 v1

@@ Abstract
Short-term Treatment Effects, Long-term Treatment Effects, Pareto Optimization, Policy Learning
I Introduction
@@
Although quite a lot works have investigated the treatment effect estimation on short-term and long-term outcomes, the problem of conflicts between them is rarely discussed. A pertinent example is the administration of medication: a higher dose may accelerate short-term recovery yet potentially cause serious long-term side effects. Moreover, the conflicts between short- and long-term outcomes have another meaning, i.e. the optimization directions for various tasks can also conflict when training estimation models. Therefore, how to trade-off between the short-term and long-term causal effects to achieve optimal treatment remains an open challenge. Last but not least, previous works mainly focus on binary treatment cases and directly regressing outcomes together without fully utilizing the information [25]. However, the nuanced scenario of continuous treatments and their impact on multiple outcomes is less explored. This gap is notable in the context of confounder balancing and policy learning. When dealing with continuous treatments, for instance, the application of balanced representation learning to estimate multiple outcomes concurrently can lead to significant information loss.
@@
  • We propose a novel Pareto-efficient algorithm containing POE and POPL, which offers not only accurate estimations of short-term and long-term outcomes but also maximal reward from effective policy learning.
@@
II Related Work
Causal inference is a powerful tool in data-related fields, enabling a deeper understanding and explanation of the complex relationships between data, particularly in recommendation systems and SQL queries [26, 27, 28, 29]. It assists in revealing the causal dynamics behind query results, offering more profound insights into the data [30, 31]. Furthermore, the significance of multi-term outcome responses and policy learning in the data management community cannot be overstated [32, 33, 34, 35]. These topics reflect the rapidly evolving landscape of data science and analytics, highlighting the field’s continuous advancement. Next, we will further discuss three directions that are highly relevant to this paper.
II-A Long-term Treatment Effect Estimation
@@
II-B Multi-task Learning
@@
II-C Policy Learning
In this work, we still focus on policy learning for Pareto-optimal policy to trade-off the short-term and long-term outcomes for maximizing the total reward. Numerous studies have been dedicated to policy learning from observational data, which can be broadly categorized into two main streams: statistics [45, 46, 47] and machine learning [48, 49, 50]. The first category primarily tackles empirical maximization problems and delves into their various relaxations, aiming to refine the theoretical underpinnings of policy learning. The second category, on the other hand, concentrates on enhancing the practical performance of policy learning techniques, with a particular emphasis on the application of doubly robust objectives. This bifurcation of focus not only delineates the diverse methods employed in policy learning but also underscores the multifaceted nature of this area. In this area, the study of adaptive paywalls in [33] explores the balance between user satisfaction and cost. Similarly, research in [34] presents a strategy to reduce revenue loss from ad blockers. Additionally, the Sim2Rec framework [35], used in sequential recommendation systems, focuses on policies that increase user engagement for long-term benefits. Together, these studies emphasize the importance of balancing different outcomes in policy learning.
III Problem Setup
III-A Notations and Assumptions
@@ III-B Preliminary
Inspired by [52], we apply continuous Pareto optimization to update the parameters of networks, and find the optimal representation in the Pareto frontier for potential outcomes estimation.
@@ IV Methodology
IV-A Pareto-Optimal Estimation
@@
IV-A1 Confounder Balancing in Continuous Treatment Cases
@@
IV-A2 Shared representation for predicting both short-term and long-term outcomes
@@ IV-A3 Pareto Optimization
IV-B Pareto-Optimal Policy Learning
In previous section, we introduced a novel Pareto-Optimal Estimation, crafted to accurately estimate potential short-term and long-term outcomes for any specific manipulated intervention. Consequently, a direct motivation emerges: to discover an optimal policy that assists practitioners in identifying the most effective treatment for maximizing the reward.
In the real-world application, the observations typically are only the pre-treatment variable covariates $X$, thus, the objective of the optimal policy is to explore which values of $T$ would lead to Pareto optimal solutions of the short-term and long-term outcomes. Similar to the Pareto estimation module described above, we train a deterministic policy backbone $\Pi(X):\mathcal{X}\rightarrow\mathcal{T}\subset\mathbb{R}$. With the estimator backbone, potential outcomes of the decided policy $\Pi(X)$ can be calculated. In order to train the policy backbone, its objective is to minimize the regret loss of each potential outcome, which is defined as the difference between the maximum outcome minus the expected outcome of $\Pi(X)$. As for the short-term outcome, the regret loss can be formally expressed as
@@ IV-B Pareto-Optimal Policy Learning
We also utilize the Pareto optimization to trade-off between the two regrets and further update this backbone. Details are described in Algorithm 2 to update the Pareto set in each step.
@@ V-A Datasets
Note that the covariates are quite different among the three semi-synthetic datasets and the simulated dataset, and we apply different equations to generate treatments and outcomes so as to guarantee the conflicts between $S$ and $Y$.
@@ V-B Baselines
TARNet. It is a classic model applied in the setting of binary treatment. There is a representation network to learn a shared embedding of $X$ across the treated group and the control group, together with two hypothesis networks to predict the outcomes of each group. This is because the dimension of $X$ is often high, while that of $T$ is only $1$. The effect of $T$ in counterfactual prediction is likely to be weakened. By separately learning for each value of $T$, its impact can be highlighted and distinguished from that of $X$.
@@ V-C2 Policy Learning
For each dataset, we randomly select two samples as representatives and plot the three-dimensional visualization of the decision-making. Specifically, we discretize the interval $\mathcal{T}$ of treatment by uniform sampling, and regard the sampled points as possible values $t^{cf}$ for counterfactual prediction. Afterwards, we iterate over all these $t^{cf}$, and use the generation formulas (mentioned in Section V-A) to yield the corresponding short-term outcome $s^{cf}$ and long-term outcome $y^{cf}$. These counterfactual outcomes are depicted as grey points in Fig. 3. The main purpose is to determine the Pareto frontier between short-term and long-term outcomes for these points, and the frontier is marked in blue. Note that larger values for both $S$ and $Y$ are preferred in our setting, meaning that the Pareto frontier is the upper-right boundary of these points.
The optimal treatment $t^{*}$ for a sample is predicted by the policy-learning module $\Pi(x_{i})$ in Section IV-B. We mark the data point $(s^{*},t^{*},y^{*})$ as a red star in Fig. 3. It can be seen that the red star is located on the Pareto frontier, meaning that the outcomes of $t^{*}$ determined by our model will not be dominated by the outcomes of any other values for $T$. The effectiveness of policy learning is validated in each dataset.
V-C3 Ablation Study
We conduct the ablation study in both the simulated and semi-synthetic datasets, whose results are demonstrated in Table V. The model called CFR-$S$ and CFR-$Y$ here correspond to the baseline CFR mentioned in the treatment effect estimation experiment. Data reported in these two rows are the same with those in Table II and Table III. We decompose it into CFR-$S$ and CFR-$Y$ primarily to facilitate a more intuitive comparison with Joint CFR. In our statement in Section V-C1, outcome $Y$ contains the information about $S$ and $X$. Therefore, it can facilitate mutual enhancement to some extent if multiple tasks are learned jointly. The improvements in both the $\mathbf{MSE_{s}}$ and $\mathbf{MSE_{y}}$ validate our viewpoint. Due to the sparsity of samples in the IHDP dataset, the joint optimization of $S$ and $Y$ leads to a slight improvement of performance. However, on the other three datasets, joint optimization $S$ and $Y$ has brought significant enhancements. Specifically, there is a $48.2\%$ improvement of $\mathbf{MSE_{s}}$ and $51.3\%$ of $\mathbf{MSE_{y}}$ in the Jobs dataset. As for Twins, the error of $S$ decreases $19.7\%$ and that of $Y$ decreases $33.3\%$. It yields even more significant improvements on the Simulation dataset, with a $76.6\%$ increase in accuracy for $S$ and $73.9\%$ for $Y$.
@@ V-C4 Hyper-parameter
VI Conclusion
In this paper, we present a novel end-to-end Pareto-Efficient framework with the purpose to determine the appropriate treatment value that would achieve Pareto optimality between short-term and long-term outcomes. Our framework consists of two key components: (1) Pareto-Optimal Estimation (POE) for predicting the potential outcomes in the short-term and long-term, and (2) Pareto-Optimal Policy Learning (POPL) for identifying the treatment value of Pareto optimality between multiple objectives. The backbones of both modules are designed and trained in a similar manner. In POE, we employ two regression losses and one representation loss to capture the prediction accuracy of the short-term and long-term potential outcomes. Meanwhile, the POPL module leverages two regret losses, which act as guiding signals to determine the Pareto optimal treatment value.

### 12 Mar 2024 v2

@@ Abstract
Short-term Treatment Effects, Long-term Treatment Effects, Pareto Optimization, Policy Learning.
1 Introduction
@@
Although many works have investigated the treatment effect estimation on short-term and long-term outcomes, the problem of conflicts between them is rarely discussed. A pertinent example is the administration of medication: a higher dose may accelerate short-term recovery yet potentially cause serious long-term side effects. Moreover, the conflicts between short- and long-term outcomes have another meaning, i.e. the optimization directions for various tasks can also conflict when training estimation models. Therefore, how to trade-off between the short-term and long-term causal effects to achieve optimal treatment remains an open challenge. Last but not least, previous works mainly focus on binary treatment cases and directly regressing outcomes together without fully utilizing the information [25]. However, the nuanced scenario of continuous treatments and their impact on multiple outcomes is less explored. This gap is notable in the context of confounder balancing and policy learning. When dealing with continuous treatments, for instance, the application of balanced representation learning to estimate multiple outcomes concurrently can lead to significant information loss.
@@
  • We propose a novel Pareto-efficient algorithm that contains two modules named POE and POPL. It can offer not only accurate estimations of short-term and long-term outcomes but also maximal reward from effective policy learning.
@@
2 Related Work
Causal inference is a powerful tool in data-related fields, enabling a deeper understanding and explanation of the complex relationships between data, such as in recommendation systems [26, 27], network embedding [28], and SQL queries [29, 30]. It assists in revealing the causal dynamics behind query results, offering more profound insights into the data [31, 32]. Furthermore, the significance of multi-term outcome responses and policy learning in the data management community cannot be overstated [33, 34, 35, 36]. These topics reflect the rapidly evolving landscape of data science and analytics, highlighting the field’s continuous advancement. Next, we will further discuss three directions that are highly relevant to this paper.
2.1 Long-term Treatment Effect Estimation
@@
2.2 Multi-task Learning
@@
2.3 Policy Learning
In this work, we still focus on policy learning for Pareto-optimal policy to trade-off between the short-term and long-term outcomes, with the purpose of maximizing the total reward. Numerous studies have been dedicated to policy learning from observational data, which can be broadly categorized into two main streams: statistics [46, 47, 48] and machine learning [49, 50, 51]. The first category primarily tackles empirical maximization problems and delves into their various relaxations, aiming to refine the theoretical underpinnings of policy learning. The second category, on the other hand, concentrates on enhancing the practical performance of policy learning techniques, with a particular emphasis on the application of doubly robust objectives. This bifurcation of focus not only delineates the diverse methods employed in policy learning but also underscores the multifaceted nature of this area. In this area, the study of adaptive paywalls in [34] explores the balance between user satisfaction and cost. Similarly, research in [35] presents a strategy to reduce revenue loss from ad blockers. Additionally, the Sim2Rec framework [36], used in sequential recommendation systems, focuses on policies that increase user engagement for long-term benefits. Together, these studies emphasize the importance of balancing different outcomes in policy learning.
3 Problem Setup
3.1 Notations and Assumptions
@@ 3.2 Preliminary
Inspired by [53], we apply a continuous Pareto optimization algorithm to update the parameters of networks, and explore the optimal representation on the Pareto frontier for potential outcomes estimation.
@@ 4 Methodology
4.1 Pareto-Optimal Estimation
@@
4.1.1 Confounder Balancing for Continuous Treatment
@@
4.1.2 Shared representation for predicting both short-term and long-term outcomes
@@ 4.1.3 Pareto Optimization
4.2 Pareto-Optimal Policy Learning
In previous section, we introduced a novel Pareto-Optimal Estimation module, crafted to accurately estimate potential short-term and long-term outcomes for any specific manipulated intervention. Consequently, a direct motivation emerges: to discover an optimal policy that assists practitioners in identifying the most effective treatment for maximizing the reward.
In the real-world application, the observations typically are only the pre-treatment variable covariates $X$, thus, the objective of the optimal policy is to explore which values of $T$ would lead to Pareto optimal solutions of the short-term and long-term outcomes. Similar to the Pareto estimation module described above, we train a deterministic policy backbone $\Pi(X):\mathcal{X}\rightarrow\mathcal{T}\subset\mathbb{R}$. With the estimator backbone, potential outcomes of the decided policy $\Pi(X)$ can be calculated. In order to train the policy backbone, its objective is to minimize the regret loss of each potential outcome, which is defined as the difference between the maximum outcome minus the expected outcome of $\Pi(X)$. As for the short-term outcome, the regret loss can be expressed as:
@@ 4.2 Pareto-Optimal Policy Learning
We also utilize the continuous Pareto optimization algorithm to trade-off between the two regrets and further update this backbone. Details are described in Algorithm 2, whose aim is to update the Pareto set in each step.
4.3 Overall
@@ 5.1 Datasets
Note that covariates are quite different among the three semi-synthetic datasets and the simulated dataset, and we apply different equations to generate treatments and outcomes so as to guarantee the conflicts between $S$ and $Y$.
@@ 5.2 Baselines
TARNet. It is a classic model applied in the binary treatment setting. There is a representation network to learn a shared embedding of $X$ across the treated group and the control group, together with two hypothesis networks to predict the outcomes of each group. This is because the dimension of $X$ is often high while that of $T$ is only $1$, thus weakening the effect of $T$ in counterfactual prediction. By separately learning for each value of $T$, its impact can be highlighted and distinguished from that of $X$.
@@ 5.3.2 Policy Learning
For each dataset, we randomly select two samples as representatives and plot the three-dimensional visualization of the decision-making. Specifically, we discretize the interval $\mathcal{T}$ of treatment by uniform sampling, and regard the sampled points as possible values $t^{cf}$ for counterfactual prediction. Afterwards, we iterate over all these $t^{cf}$, and use the generation formulas (mentioned in Section 5.1) to yield the corresponding short-term outcome $s^{cf}$ and long-term outcome $y^{cf}$. These counterfactual outcomes are depicted as grey points in Fig. 3. The main purpose is to determine the Pareto frontier between short-term and long-term outcomes for these points, and the frontier is marked in blue. Note that larger values for both $S$ and $Y$ are preferred in our setting, meaning that the Pareto frontier is the upper-right boundary of these points.
The optimal treatment $t^{*}$ for a sample is predicted by the policy-learning module $\Pi(x_{i})$ in Section 4.2. We mark the data point $(s^{*},t^{*},y^{*})$ as a red star in Fig. 3. It can be seen that the red star is located on the Pareto frontier, meaning that the outcomes of $t^{*}$ determined by our model will not be dominated by the outcomes of any other values for $T$. The effectiveness of policy learning is validated in each dataset.
5.3.3 Ablation Study
We conduct the ablation study in both the simulated and semi-synthetic datasets, whose results are demonstrated in Table V. The model called CFR-$S$ and CFR-$Y$ here correspond to the baseline CFR mentioned in the treatment effect estimation experiment. Data reported in these two rows are the same with those in Table II and Table III. We decompose it into CFR-$S$ and CFR-$Y$ primarily to facilitate a more intuitive comparison with Joint CFR. In our statement in Section 5.3.1, outcome $Y$ contains the information about $S$ and $X$. Therefore, it can facilitate mutual enhancement to some extent if multiple tasks are learned jointly. The improvements in both the $\mathbf{MSE_{s}}$ and $\mathbf{MSE_{y}}$ validate our viewpoint. Due to the sparsity of samples in the IHDP dataset, the joint optimization of $S$ and $Y$ leads to a slight improvement of performance. However, on the other three datasets, joint optimization $S$ and $Y$ has brought significant enhancements. Specifically, there is a $48.2\%$ improvement of $\mathbf{MSE_{s}}$ and $51.3\%$ of $\mathbf{MSE_{y}}$ in the Jobs dataset. As for Twins, the error of $S$ decreases $19.7\%$ and that of $Y$ decreases $33.3\%$. It yields even more significant improvements on the Simulation dataset, with a $76.6\%$ increase in accuracy for $S$ and $73.9\%$ for $Y$.
@@ 5.3.4 Hyper-parameter
6 Conclusion
In this paper, we propose a novel end-to-end Pareto-Efficient framework with the purpose to determine the appropriate treatment value that would achieve Pareto optimality between short-term and long-term outcomes. Our framework consists of two key components: (1) Pareto-Optimal Estimation (POE) for predicting the potential outcomes in the short-term and long-term, and (2) Pareto-Optimal Policy Learning (POPL) for identifying the treatment value of Pareto optimality between multiple objectives. The backbones of both modules are designed and trained in a similar manner. In POE, we employ two regression losses and one representation loss to capture the prediction accuracy of the short-term and long-term potential outcomes. The POPL module leverages two regret losses, which act as guiding signals to determine the Pareto optimal treatment value.